OpenStudy (anonymous):

There are 52 white keys and 36 black keys on a standard piano. What is the probability that if a person did not look and hit the keyboard twice that he or she would first hit a white key and then hit a black key?

OpenStudy (anonymous):

\[\frac{52}{88}\times \frac{36}{88}\] maybe can reduce

OpenStudy (kinggeorge):

There are a total of \(52+36=88\) keys on the piano. Thus, the probability of hitting a white key would be \(52 \over 88\) and a black key \(36 \over 88\) Therefore, the probability of hitting a white key then a black key is given by \[{52 \over 88} \times {36 \over 88}\]

OpenStudy (anonymous):

thank you for explaining it too me :) its easier to understand now how to do the problem

OpenStudy (kinggeorge):

you're welcome

OpenStudy (anonymous):

So the answer is 1/1872 probablity?

OpenStudy (anonymous):

Could someone help me with this , because I got 117/484.

OpenStudy (kendricklamar2014):

\[\frac{ 52 }{ 88 } \times \frac{ 36 }{ 88 } = \frac{ 117 }{ 484 }\]

OpenStudy (kendricklamar2014):

Just clarifying the answer ;)